Evaluate the arithmetic series where t1=42, tn=146, and n=14.

Answer by jim_thompson5910(35256)

Sn = (n*(t1+tn))/2

S14 = (14*(42+146))/2

S14 = (14*(188))/2

S14 = 2632/2

S14 = 1316

So the 14th partial sum is 1316